OpenStudy (anonymous):

What is the magnitude and direction of if G(2, -2) and H(7, 6)?

OpenStudy (anonymous):

????????????????

OpenStudy (anonymous):

Magnitude is \(\sqrt{{(x_0-x_1)}^2+{(y_0-y_1)}^2} = \sqrt{{(2-7)}^2+{(-2-6)}^2} = \sqrt{25+64} = \sqrt{89}\).

OpenStudy (fifciol):

magnitude: \[m=\sqrt{(7-2)^2+(6-(-2))^2}=9,43\] \[\tan \theta=\frac{ 6-(-2) }{7-2 } \rightarrow \theta=58^{0}\]

OpenStudy (anonymous):

The direction could also be 360-58=302°, since we don't know if it is \(\vec{GH}\) or \(\vec{HG}\)

OpenStudy (fifciol):

yes, it could be 58 or -58

OpenStudy (anonymous):

Oh, wait sorry, not 360-58=302, I meant 180+58=238